A mail-order house uses 16,260 boxes a year. Carrying costs are 60 cents per box a year, and ordering costs are $96. The following price schedule applies.

Number of Boxes Price per Box
1,000 to 1,999 $1.25                
2,000 to 4,999 1.20                
5,000 to 9,999 1.15                
10,000 or more 1.10                
a.

Determine the optimal order quantity. (Round your answer to the nearest whole number.)

  Optimal order quantity boxes
b.

Determine the number of orders per year. (Round your answer to 2 decimal places.)

  Number of order per year  
Answer #1

DEMAND = 16260

HOLDING COST = 0.6

ORDERING COST = 96


EOQ = SQRT(2DS/H)

Optimal order quantity = Quantity with lowest total annual cost of inventory

TCI = annual holding cost + annual ordering cost + annual material cost

Annual holding = adjusted quantity / 2 * holding cost

Annual ordering = demand / adjusted quantity * ordering cost

Annual material cost = demand * cost per unit


ADJUSTED Q = EOQ IF EOQ > LOWER LIMIT, < UPPER LIMIT

LOWER LIMIT IF EOQ < LOWER LIMIT

UPPER LIMIT IF EOQ > UPPER LIMIT



NO.

LOWER BRACKET

UPPER BRACKET

PER UNIT

HC

EOQ

Q*

TOTAL COST

FORMULA

1

1

1999

1.25

0.6

2281

1999

21705.57

(16260 * 1.25) + ((1999 / 2) * 0.6) + ((16260 / 1999) * 96) = 21706

2

2000

4999

1.2

0.6

2281

2281

20880.63143

(16260 * 1.2) + ((2281 / 2) * 0.6) + ((16260 / 2281) * 96) = 20881

3

5000

9999

1.15

0.6

2281

5000

20511.19

(16260 * 1.15) + ((5000 / 2) * 0.6) + ((16260 / 5000) * 96) = 20511

4

10000

MORE

1.1

0.6

2281

10000

21042.1

(16260 * 1.1) + ((10000 / 2) * 0.6) + ((16260 / 10000) * 96) = 21042


OPTIMAL ORDER QUANTITY = 5000

TOTAL COST OF INVENTORY = 20511



2. NO. OF ORDERS PER YEAR = D / Q = 16260 / 5000 = 3.25



LEAVE A LIKE IF YOU THINK THIS ANSWER WAS HELPFUL.